# Adaptive Off-Policy Inference for M-Estimators Under Model Misspecification

### v1

@@ Abstract
When data are collected adaptively, such as in bandit algorithms, classical statistical approaches such as ordinary least squares and $M$-estimation will often fail to achieve asymptotic normality. Although recent lines of work have modified the classical approaches to ensure valid inference on adaptively collected data, most of these works assume that the model is correctly specified. We propose a method that provides valid inference for M-estimators that use adaptively collected bandit data with a (possibly) misspecified working model. A key ingredient in our approach is the use of flexible machine learning approaches to stabilize the variance induced by adaptive data collection. A major novelty is that our procedure enables the construction of valid confidence sets even in settings where treatment policies are unstable and non-converging, such as when there is no unique optimal arm and standard bandit algorithms are used. Empirical results on semi-synthetic datasets constructed from the Osteoarthritis Initiative demonstrate that the method maintains type I error control, while existing methods for inference in adaptive settings do not cover in the misspecified case.
@@ 1 Introduction
Adaptive data collection has become a common practice in modern data analysis due to the rise of reinforcement learning and sequential data collections in contexts such as personalized healthcare (Yom-Tov et al., 2017), web recommendations (Afsar et al., 2022), and clinical trials (Zhao et al., 2009; Liu et al., 2020). In contrast to classical settings, where i.i.d. data are observed, adaptively collected data allows an analyst to interact with the decision-making algorithm in various ways, leading to sequentially dependent data. A canonical example is the (contextual) bandit problem where an analyst is allowed to make a choice of treatment (or “arm”) at each time step and then observes a reward. Bandit algorithms attempt to maximize the total expected reward over time. These algorithms tend to result in data collection policies that explore treatment options in early rounds, and they tend to shift to a greedier strategy that chooses the treatment with the highest estimated reward in the later rounds. When inference is conducted on data collected by bandit algorithms, the greedy search process can result in unstable variances (Deshpande et al., 2018; Zhang et al., 2020) and biased estimates of the average reward in each treatment arm (Shin et al., 2019) that make naive approaches to inference invalid.
There is a rich literature to adjust to the difficulties that arise for inference on adaptively collected data. Typical approaches focus on limiting the variance in each data collection using techniques such as propensity score truncation (Cook et al., 2024) or post-hoc reweighting of the data when constructing an estimator (Bibaut et al., 2021; Syrgkanis and Zhan, 2024). These approaches allow for assumption-lean inference in semiparametric settings, but they limit the study to relatively simple functionals such as the average reward for a fixed treatment or contrasts between average rewards for fixed arms.
An alternative set of approaches assumes a linear model linking treatment ($X_{t}$) and rewards ($Y_{t}$),
@@
where $\epsilon_{t}$ is i.i.d. random noise. It is striking that even in this simple setting, estimators such as ordinary least squares fail to achieve asymptotic normality when $X_{t}$ depends on previous rounds $\{X_{i},Y_{i}\}_{i=1}^{t-1}$. A classical result (Lai and Wei, 1982) demonstrates that the OLS estimate for $\theta$ is asymptotically normal when the covariance matrix $(\sum_{t=1}^{T}X_{t}X_{t}^{T})^{-1}$ converges to a fixed deterministic sequence. More recent literature (Zhang et al., 2020; Deshpande et al., 2018; Khamaru et al., 2021) makes the point that this assumption typically does not apply for common bandit algorithms when the margin (that is, the difference in mean rewards between arms) is zero. Zhang et al. (2020) proposes a solution through batching, by fixing the sampling rule for each batch and letting the number of observations within each batch go to infinity. This means that even if the sampling rule does not concentrate across time, it is fixed within each batch, allowing the empirical covariance matrix within each batch to concentrate. An alternative set of approaches (Deshpande et al., 2018; Khamaru et al., 2021) uses an online debiasing approach, where the finite sample bias is controlled through regularization. Alternative approaches use concentration results to construct confidence regions that are finite-sample valid uniformly over time and using estimation equations (Ying et al., 2023).
@@
$\theta^{\star}:=\operatorname*{argmax}_{\theta}\mathbb{E}\left[\left(Y_{t}-\theta^{T}X_{t}\right)^{2}\right].$
@@
Moving beyond the linear case, Zhang et al. (2021) considers statistical inference on $M$-estimators in the contextual bandit problem, allowing for more complex models to describe the reward structure, such as generalized linear models. However, this work again requires that the conditional mean of the model is correctly specified in the working model. A related paper (Zhang et al., 2023) performs inference for $Z$-estimators and allows for a misspecified model, but only under a finite amount of adaptivity. In this framework, $n$ individuals are tracked over $T$ time periods, with $T$ fixed and $n$ growing to $\infty$. The advantage of this approach is that it allows for non-stationary reward and context distributions over time, but it relies on the number of individuals enrolled in the trial to tend to $\infty$ and does not allow adaptive decisions to be made separately for each individual in the trial.
@@ 1.1 Our Contributions
In this work, we propose a methodology for inferring parameters in $M$-estimation problems under a potentially misspecified working model and non-finite levels of adaptivity. The major result, Theorem 1, provides a Central Limit Theorem that enables the construction of confidence intervals when the variance (which varies over time and may not converge) of the score function can be estimated consistently.
The problem of estimating the variance of the score function is nontrivial because using naive empirical estimates when applying Theorem 1 is only valid when the variance is a fixed deterministic quantity. In many adaptive settings, such as bandit problems where expected rewards between arms are comparable, action selection probabilities will be non-convergent random quantities, leading to unstable variances (Zhang et al., 2020). To solve this problem, we leverage flexible machine learning methods to construct valid time-varying plug-in estimates for the variances at each time step.
In a work that is concurrent to ours (the authors having released their work one week prior to the first version of this manuscript), Guo and Xu (2025) also provides a procedure for inference on $Z$-estimators trained on adaptively collected data via inverse propensity weighting, but only under the assumption that treatment policies converge as $T\rightarrow\infty$. Since many common bandit algorithms fail to converge under model misspecification, the authors provide a sufficient set of conditions to guarantee policy convergence (e.g. continuous policies that are sufficiently smooth). We take a different approach — rather than restricting ourselves to only particular classes of policy that are guaranteed to converge, we allow the analyst to assign treatments in relatively arbitrary ways and then adjust for the adaptivity post hoc by estimating the conditional variance separately at each time step. In cases where the policy converges, these machine-learning-based approaches can be replaced with simple empirical estimates of the variance, and our procedure can be simplified considerably.
@@ 1.2 Paper Outline
In Section 2, we define the problem, most crucially introducing definitions of target parameters inspired by the off-policy evaluation literature that remain tractable in the misspecified setting. In Section 3, we present a CLT that enables inference for $M$-estimators in the adaptive setting given accurate (time-varying) estimates of the variance of the score function. Section 4 discusses practical strategies for estimating variance using flexible machine learning approaches. Section 5 presents empirical results on semi-synthetic datasets constructed from the Osteoarthritis Initiative, a publicly available longitudinal dataset provided by the NIH which tracks health outcomes of patients with osteoarthritis. The results verify the validity of the procedure and the failure of existing approaches to provide confidence sets with valid coverage. We provide concluding thoughts in Section 6, including potential limitations of this approach and possible directions for future work.
@@ 1.3 Notation
We introduce some shorthand notation used throughout the paper. We define $n]:=\{1,...,n\}$ for a positive integer $n$. We denote $e_{j}$ as the $j$-th standard basis vector in $\mathbb{R}^{d}$. For a function $f_{\theta}:\mathbb{R}^{d}\rightarrow\mathbb{R}$, we refer to $\dot{f}_{\theta}$ as the first derivative of $f$ with respect to $\theta$, $\ddot{f}_{\theta}$ as the Hessian matrix with respect to $\theta$, $\dddot{f}_{\theta}$ as the third derivative with respect to $\theta$, and so on. For $M\in\mathbb{R}^{d_{1}\times d_{2}\times d_{3}}$, we write $\left\lVert M\right\rVert_{1}=\sum_{i=1}^{d_{1}}\sum_{i=1}^{d_{2}}\sum_{i=1}^{d_{3}}|a_{i,j,k}|$. For two matrices $A,B$, $A\succeq B$ means that ABA-B is positive semidefinite.
@@ 2 Problem Setup
We assume a stochastic bandit environment in which features $X_{t}\in\mathbb{R}$ are observed in each round $t\in[T]$. After observing the feature, an analyst chooses an action $A_{t}$ from a set of actions $\mathcal{A}$. For simplicity, we assume $\mathcal{A}:=\{1,...,K\}$ is a finite set of $K$ possible discrete actions. After selecting an action, an outcome $Y_{t}$ is observed. We denote $\{(Y_{t}(a)):a\in\mathcal{A}\}$ to be the counterfactual outcome had the analyst chosen $a$ in round $t$, regardless of the actual action taken. We assume that the joint distribution of features and potential outcomes is independent and identically distributed across time.
@@
Although potential outcomes are distributed i.i.d., we allow the analyst to choose the treatment adaptively based on a pooled history $\mathcal{H}_{t-1}:=\{(X_{i},A_{i},Y_{i})\}_{i=1}^{t-1}$. Formally, we say $\pi_{t}(a|X_{t}):=\mathbb{P}(A_{t}=a|X_{t},\mathcal{H}_{t-1})\in\sigma(\mathcal{H}_{t-1})$ for all $a\in\mathcal{A}$. We can summarize the overall joint density over actions, features, and outcomes at time step $t$ as
@@
Our goal is to construct confidence regions for a parameter $\theta^{\star}$, which we define as the expected maximizer of some function $m_{\theta}$ in some parameter space $\Theta$. For example, $m_{\theta}(x,a,y):=(y-\theta^{T}(x,a))^{2}$ corresponds to an ordinary least squares regression, and $m_{\theta}(x,a,y):=-y(x,a)^{T}\theta+\psi((x,a)^{T}\theta)$, where $\psi$ denotes the convex log-partition function, corresponds more generally to GLMs (Zrnic and Candès, 2024). Some nuance is in order to determine the distribution over which the expected loss is minimized. The most natural choice, following Equation 1, is to define $\theta^{\star}$ as
$\operatorname*{argmax}_{\theta\in\Theta}\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}m_{\theta}(X_{t},A_{t},Y_{t})\right]=\operatorname*{argmax}_{\theta\in\Theta}\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}_{\mathcal{P},\pi_{t}}\left[m_{\theta}(X_{t},A_{t},Y_{t})\mid\mathcal{H}_{t-1}\right]\right],$
Unfortunately, this is often not a tractable object of inference due to the adaptivity of $\pi_{t}$. Instead, we tackle the problem of off-policy learning and seek inference under a hypothetical policy $\pi_{e}(a|x)$ that is distinct from the policy used during treatment and invariant across time.
@@
The choice of $\pi_{e}(a|x)$ may be a fixed and known distribution of interest. For example, $p_{e}(A_{t}=a|X_{t})=1/|\mathcal{A}|\text{ for all }a\in\mathcal{A}$ and $t\in[T]$ would represent a hypothetical setting in which treatments are chosen uniformly within the population. Alternatively, $\pi_{e}(a|x)$ might be a fixed but unknown quantity to be estimated later, such as a hypothetical optimal policy or a limiting quantity such as $\lim_{t\rightarrow\infty}\mathbb{P}(A_{t}=a|X_{t},\mathcal{H}_{t-1})$ (which we do not assume to exist in general in our paper). An appeal to Slutsky’s theorem will allow us to substitute estimated versions of optimal policies after data collection to form our estimators, so long as the estimates do not depend on the history as $T\rightarrow\infty$. Of course, the interpretation of $\theta^{\star}$ can vary dramatically depending on the choice of $\pi_{e}$ as a simple example will elucidate.
Figure 1: Illustration for Example 1. Under the misspecified linear model $m_{\theta}=(Y_{t}-\theta_{0}-\theta_{1}A_{t})^{2}$, $(\theta_{0}^{\star},\theta_{1}^{\star})=(-0.2,4)$ for the first policy and $(\theta_{0}^{\star},\theta_{1}^{\star})=(-5.3,15)$ for the second policy. Intuitively, the misspecified model will try to estimate a secant line at different points of the quadratic function, resulting in very different interpretations for the target parameter. As such, the choice of evaluation policy is critical for properly interpreting the target parameter.
@@
Consider data generated as $Y_{t}\sim N(6A_{t}^{2},1)$ under two treatment regimes, where $\mathcal{A}=\{0,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1\}$. The first policy is uniform over $A_{t}\in\{0,0.1,0.2,0.3,0.4,0.5\}$. The second policy is uniform over $A_{t}\in\{0.6,0.7,0.8,0.9,1.0\}$. Under the correctly specified model $m_{\theta}(A_{t},Y_{t}):=(Y_{t}-\theta_{0}-\theta_{1}A_{t}^{2})^{2}$, $(\theta_{0}^{\star},\theta_{1}^{\star})=(0,6)$ for both policies. On the other hand, under an incorrectly specified linear model, $\theta_{0}^{\star}$ and $\theta_{1}^{\star}$ can vary markedly as seen in Figure 1.
@@
We can use this framework to target the average treatment effect by using a vector of indicator variables as the chosen model with a square loss function. We can permit any evaluation policy that satisfies the standard causal assumption of $Y_{t}(a)\perp A_{t}$ under $\pi_{e}$, and such that $\pi_{e}(A_{t}=a)>0$.
Formally, we let $\theta=(\theta_{1},...,\theta_{|\mathcal{A}|})$ and consider setting $m_{\theta}=(Y_{t}-\sum_{a\in\mathcal{A}}1_{A_{t}=a}\theta_{a})^{2}$. First order conditions imply that for all $t$,
$0=\mathbb{E}_{\mathcal{P},\pi_{e}}\left[Y_{t}-\sum_{a\in\mathcal{A}}1_{A_{t}=a}\theta^{\star}_{a}\right].$
Rearranging terms gives us that
@@
To infer the average effect of treatment $j$ versus treatment $k$, we can then target $\eta^{T}\theta^{\star}$ for some contrast vector $\eta:=e_{j}-e_{k}$. For example (1,1)(-1,1) in the binary setting would recover the typical formulation of the average treatment effect
$\mathbb{E}\left[Y(1)-Y(0)\right].$
@@
Following Zhang et al. (2021), one option is to consider estimators of the form
@@
When this is true conditionally, it will also be true marginally for any choice of policy, therefore under no model misspecification, we have at every time step $t$ that,
@@
eliminating the need to consider the choice of policy when defining the target parameter. This also means that any weights $w_{t}\in\sigma(X_{t},A_{t},\mathcal{H}_{t-1})$ can be used without affecting the first moment of the estimator $\hat{\theta}_{0}$. We can see this because Equation 3 implies that when the score function is set equal to 0, we have that
@@
for all $X_{t},A_{t},w_{t}\in\sigma(X_{t},A_{t},\mathcal{H}_{t-1})$. Thus, $w_{t}$ can be arbitrarily chosen based on $X_{t}$, $A_{t}$, and $\mathcal{H}_{t-1}$ without violating the first-order conditions. For this reason, previous work often uses $w_{t}$ as a free parameter to control the variance without having to worry that this type of re-weighting will bias $\hat{\theta}_{T}$. As an example, Zhang et al. (2021) chooses $w_{t}=\sqrt{\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}(A_{t}|X_{t},\mathcal{H}_{t-1})}}$, which allows the variance of the score function to converge to a value independent of history. Other examples include online debiasing approaches (Deshpande et al., 2018; Khamaru et al., 2021) which use regularization to choose $w_{t}$ so that the variance of $\hat{\theta}_{T}$ is controlled.
@@
which is only for the specific choice of $w_{t}=\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}(A_{t}|X_{t},\mathcal{H}_{t-1})}$. Thus, we are constrained to use inverse propensity weighting to define an estimator such as
@@
Using inverse propensity weights ensures that $\hat{\theta}_{0}$ will be unbiased for $\theta^{\star}$, but we must now consider alternative methods to stabilize the variance of the score function, which we will discuss in the next section.
@@
The estimator potentially improves power when the model $f_{t}$ is well trained by augmenting observations with predictions from a model. This approach is routinely used when estimating simple functionals like average treatment effect in causal inference (Chernozhukov et al., 2018) and off-policy evaluation (Hadad et al., 2021) and is referred to as augmented inverse probability weighted estimator. More recently, Zrnic and Candès (2024) uses $M$-estimators with augmented observations from machine learning to improve statistical efficiency but only in the non-causal, non-adaptive observational setting.

### v2

@@ Abstract
When data are collected adaptively, such as in bandit algorithms, classical statistical approaches such as ordinary least squares and $M$-estimation will often fail to achieve asymptotic normality. Although recent lines of work have modified the classical approaches to ensure valid inference on adaptively collected data, most of these works assume that the model is correctly specified. The misspecified setting poses unique challenges because the parameter of interest itself may not be well-defined over a non-stationary distribution of rewards. We therefore tackle the problem of off-policy inference in adaptive settings, where we uniquely define a projected solution over a stationary evaluation policy. Our method provides valid inference for $M$-estimators that use adaptively collected bandit data with a possibly misspecified working model. A key ingredient in our approach is the use of flexible approaches to stabilize the variance induced by adaptive data collection. A major novelty is that the procedure enables the construction of valid confidence sets even in settings where treatment policies are unstable and non-converging, such as when there is no unique optimal arm and standard bandit algorithms are used. Empirical results on semi-synthetic datasets constructed from the Osteoarthritis Initiative demonstrate that the method maintains type I error control, while existing methods for inference in adaptive settings do not cover in the misspecified case.
@@ 1 Introduction
Adaptive data collection has become a common practice in modern data analysis due to the rise of reinforcement learning and sequential data collections in contexts such as personalized healthcare (Yom-Tov et al., 2017), web recommendations (Afsar et al., 2022), and clinical trials (Zhao et al., 2009; Liu et al., 2020). In contrast to classical settings, where i.i.d. data are observed, adaptively collected data allows an analyst to interact with the decision-making algorithm in various ways, leading to sequentially dependent data. A canonical example is the (contextual) bandit problem where an analyst is allowed to make a choice of treatment (or “arm”) at each time step and then observes a reward. Bandit algorithms attempt to maximize the total expected reward over time. These algorithms tend to result in data collection policies that explore treatment options in the early rounds and tend to shift to a greedier strategy that chooses the treatment with the highest estimated reward in the later rounds. When inference is conducted on data collected by bandit algorithms, the greedy search process can result in unstable variances (Deshpande et al., 2018; Zhang et al., 2020) and biased estimates of the average reward in each treatment arm (Shin et al., 2019) that make naive approaches to inference invalid.
There is a rich literature for adapting to the difficulties that arise for inference on adaptively collected data. Typical approaches focus on limiting the variance in each data collection using techniques such as propensity score truncation (Cook et al., 2024) or post-hoc reweighting of the data when constructing an estimator (Bibaut et al., 2021; Syrgkanis and Zhan, 2024). These approaches allow for assumption-lean inference in semiparametric settings, but they limit the study to relatively simple functionals such as the average reward for a fixed treatment or contrasts between average rewards for fixed arms.
An alternative set of approaches assumes a linear model linking treatment ($X_{t}\in\mathbb{R}^{d}$) and rewards ($Y_{t}\in\mathbb{R}$),
@@
where $\epsilon_{t}$ is i.i.d. random noise. It is striking that even in this simple setting, estimators such as ordinary least squares fail to achieve asymptotic normality when $X_{t}$ depends on previous rounds $\{X_{i},Y_{i}\}_{i=1}^{t-1}$. A classical result (Lai and Wei, 1982) demonstrates that a necessary condition for the OLS estimate of $\theta$ to be asymptotically normal is when there exists a deterministic sequence of positive definite matrices $\{B_{T}\}_{T=1}^{\infty}$ such that $B_{T}^{-1}\sum_{t=1}^{T}X_{t}X_{t}^{T}\overset{p}{\to}I_{d}$. More recent literature (Zhang et al., 2020; Deshpande et al., 2018; Khamaru et al., 2025) makes the point that this assumption typically does not apply for common bandit algorithms when the margin (that is, the difference in mean rewards between arms) is zero. Zhang et al. (2020) proposes a solution through batching, by fixing the sampling rule for each batch and letting the number of observations within each batch go to infinity. This means that even if the sampling rule does not concentrate across time, it is fixed within each batch, allowing the empirical covariance matrix within each batch to concentrate. An alternative set of approaches (Deshpande et al., 2018; Khamaru et al., 2025) uses an online debiasing approach, where the finite sample bias is controlled through regularization. Other approaches generalize this this to the generalized partial linear model Lin et al. (2025) and for estimating equations (Ying et al., 2023).
@@
$\theta^{\star}:=\operatorname*{argmin}_{\theta}\mathbb{E}\left[\left(Y-\theta^{T}X\right)^{2}\right]$ (1)
@@
Moving beyond the linear case, Zhang et al. (2021) considers statistical inference on $M$-estimators in the contextual bandit problem, allowing for more complex models to describe the reward structure, such as generalized linear models. However, this work again requires that the conditional mean of the model is correctly specified in the working model. However, in adaptive settings, selecting a target corresponding to a projected solution analogous to Equation 1 is not straightforward because the data-generating mechanism is non-stationary over time and dependent on user choices. We therefore define a target with respect to a fixed evaluation policy, which yields an off-policy projection parameter as an inferential target that remains interpretable under persistent adaptivity; formal details are deferred to Section 2.
An alternative approach (Zhang et al., 2023) performs inference for $Z$-estimators and allows for a misspecified model, but only under a finite amount of adaptivity. In this framework, $n$ individuals are tracked over $T$ time periods, with $T$ fixed and $n$ growing to $\infty$. The advantage of this approach is that it allows for non-stationary reward and context distributions over time, but it relies on the number of individuals enrolled in the trial to tend to $\infty$ and does not allow adaptive decisions to be made separately for each individual in the trial.
@@ 1.1 Our Contributions
In this work, we propose a methodology for inferring parameters in $M$-estimation problems under a potentially misspecified working model and non-vanishing levels of adaptivity (i.e. without assuming the adaptivity eventually vanishes). The main result, Theorem 1, provides a Central Limit Theorem that enables the construction of confidence intervals when the variance (which varies over time and may not converge) of the score function can be estimated consistently.
The problem of estimating the variance of the score function is nontrivial because using naive empirical estimates when applying Theorem 1 is only valid when the variance is a fixed deterministic quantity. In many adaptive settings, such as bandit problems where expected rewards between arms are comparable, action selection probabilities will be non-convergent random quantities, leading to unstable variances (Zhang et al., 2020). To solve this problem, we propose flexible methods to construct valid time-varying plug-in estimates for the variances at each time step.
In a work that is concurrent to ours, Guo and Xu (2025) also provides a procedure for inference on $Z$-estimators trained on adaptively collected data via inverse propensity weighting, but only under the assumption that treatment policies converge as $T\rightarrow\infty$. Since many common bandit algorithms fail to converge under model misspecification, the authors provide a sufficient set of conditions to guarantee policy convergence (e.g., continuous policies that are sufficiently smooth). We take a different approach — rather than restricting ourselves to only particular classes of policy that are guaranteed to converge, we allow the analyst to assign treatments in relatively arbitrary ways and then adjust for the adaptivity post hoc by estimating the conditional variance separately at each time step. In cases where the policy converges, these machine-learning-based approaches can be replaced with simple empirical estimates of the variance, and our procedure can be simplified considerably.
@@ 1.2 Paper Outline
In Section 2, we define the problem, most crucially introducing definitions of target parameters inspired by the off-policy evaluation literature that remain tractable in the misspecified setting. In Section 3, we present a CLT that enables inference for $M$-estimators in the adaptive setting given accurate (time-varying) estimates of the variance of the score function. Section 4 discusses practical strategies for estimating variance using flexible machine learning approaches. Section 5 presents empirical results on semi-synthetic datasets constructed from the Osteoarthritis Initiative, a publicly available longitudinal dataset provided by the NIH which tracks health outcomes of patients with osteoarthritis. The results verify the validity of the procedure and the failure of existing approaches to provide confidence sets with valid coverage. We provide concluding thoughts in Section 6, including potential limitations of this approach and possible directions for future work. All proofs for stated theorems are included in the Appendix.
@@ 1.3 Notation
We introduce some shorthand notation used throughout the paper. We define $n]:=\{1,...,n\}$ for a positive integer $n$. We denote $e_{j}$ as the $j$-th standard basis vector in $\mathbb{R}^{d}$. For a function $f_{\theta}:\mathbb{R}^{d}\rightarrow\mathbb{R}$, we refer to $\dot{f}_{\theta}$ as the first derivative of $f$ with respect to $\theta$, $\ddot{f}_{\theta}$ as the Hessian matrix with respect to $\theta$, $\dddot{f}_{\theta}$ as the third derivative with respect to $\theta$, and so on. For $M\in\mathbb{R}^{d_{1}\times d_{2}\times d_{3}}$, we write $\left\lVert M\right\rVert_{1}=\sum_{i=1}^{d_{1}}\sum_{j=1}^{d_{2}}\sum_{k=1}^{d_{3}}|a_{i,j,k}|$. For two matrices $A,B$, $A\succeq B$ means that ABA-B is positive semidefinite.
@@ 2 Problem Setup
We assume a stochastic bandit environment in which features $X_{t}\in\mathbb{R}$ are observed in each round $t\in[T]$. After observing a feature, an analyst chooses an action $A_{t}$ from a finite set of $K$ actions $\mathcal{A}:=\{1,...,K\}$. After selecting an action, an outcome $Y_{t}$ is observed. We denote $Y_{t}(a)$, where $a\in\mathcal{A}$, to be the counterfactual result had the analyst chosen $a$ in round $t$, regardless of the actual action taken. We assume that the joint distribution of features and potential outcomes is independent and identically distributed across time.
@@
Although potential outcomes are distributed i.i.d., we allow the analyst to choose the treatment adaptively based on a pooled history $\mathcal{H}_{t-1}:=\{(X_{i},A_{i},Y_{i})\}_{i=1}^{t-1}$. Formally, we say $\pi_{t}(a|X_{t}):=\mathbb{P}(A_{t}=a|X_{t},\mathcal{H}_{t-1})\in\sigma(\mathcal{H}_{t-1})$ for all $a\in\mathcal{A}$. We can summarize the conditional density (given $\mathcal{H}_{t-1}$) over actions, features, and outcomes at time step $t$ as
@@
Our goal is to construct confidence regions for a parameter $\theta^{\star}$, which we define as the expected maximizer of some function $m_{\theta}$ in some parameter space $\Theta$. For example, $m_{\theta}(x,a,y):=-(y-\theta^{T}(x,a))^{2}$ corresponds to an ordinary least squares regression, and $m_{\theta}(x,a,y):=-y(x,a)^{T}\theta+\psi((x,a)^{T}\theta)$, where $\psi$ denotes the convex log-partition function, corresponds more generally to GLMs (Agresti, 2015). Some nuance is in order to determine the distribution over which the expected loss is minimized. Because the distribution of actions evolves over time, the distribution at each time step may be substantially different so the solution to $\operatorname*{argmax}_{\theta\in\Theta}\mathbb{E}\left[m_{\theta}(X_{t},A_{t},Y_{t})\right]$ will also vary substantially across $t\in[T]$. In order to ensure there is a unique and stable inferential target, we tackle the problem of off-policy learning and seek inference under a hypothetical policy $\pi_{e}(a|x)$ that is distinct from the policy used during treatment and invariant over time.
@@
The policy $\pi_{e}(a|x)$ may be a fixed and known distribution of interest. For example, $\pi_{e}(A_{t}=a|X_{t})=1/|\mathcal{A}|\text{ for all }a\in\mathcal{A}$ and $t\in[T]$ would represent a hypothetical setting in which treatments are chosen uniformly within the population. Alternatively, $\pi_{e}(a|x)$ might be a fixed but unknown quantity to be estimated later, such as a hypothetical optimal policy or a limiting quantity such as $\lim_{t\rightarrow\infty}\mathbb{P}(A_{t}=a|X_{t},\mathcal{H}_{t-1})$ (which we do not assume to exist in general in our paper). An appeal to Slutsky’s theorem will allow us to substitute estimated versions of optimal policies after data collection to form our estimators, so long as the estimates are asymptotically independent of the history as $T\rightarrow\infty$. Of course, the interpretation of $\theta^{\star}$ can vary dramatically depending on the choice of $\pi_{e}$, as a simple example will elucidate.
Figure 1: Illustration for Example 1. Under the misspecified linear model $m_{\theta}=-(Y_{t}-\theta_{0}-\theta_{1}A_{t})^{2}$, $(\theta_{0}^{\star},\theta_{1}^{\star})=(-0.2,4)$ for the first policy and $(\theta_{0}^{\star},\theta_{1}^{\star})=(-5.3,15)$ for the second policy. Intuitively, the misspecified model will try to estimate a secant line at different points of the quadratic function, resulting in very different interpretations for the target parameter. As such, the choice of evaluation policy is critical for properly interpreting the target parameter.
@@
Consider data generated as $Y_{t}\sim N(6A_{t}^{2},1)$ under two treatment regimes, where $\mathcal{A}=\{0,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1\}$. The first policy is uniform over $A_{t}\in\{0,0.1,0.2,0.3,0.4,0.5\}$. The second policy is uniform over $A_{t}\in\{0.6,0.7,0.8,0.9,1.0\}$. Under the correctly specified model $m_{\theta}(A_{t},Y_{t}):=-(Y_{t}-\theta_{0}-\theta_{1}A_{t}^{2})^{2}$, $(\theta_{0}^{\star},\theta_{1}^{\star})=(0,6)$ for both policies. On the other hand, under an incorrectly specified linear model, $\theta_{0}^{\star}$ and $\theta_{1}^{\star}$ can vary markedly as seen in Figure 1.
@@
We can use this framework to target the average treatment effect by using a vector of indicator variables as the chosen model with a square loss function. We can permit any evaluation policy that satisfies the standard causal assumption of $Y_{t}(a)\perp\!\!\!\perp A_{t}$ under $\pi_{e}$, and such that $\pi_{e}(A_{t}=a)>0$.
Formally, we let $\theta=(\theta_{1},...,\theta_{|\mathcal{A}|})$ and consider setting $m_{\theta}=-(Y_{t}-\sum_{a=1}^{K}1_{A_{t}=a}\theta_{a})^{2}$. First order conditions imply that for all $t$,
$0=\mathbb{E}_{\mathcal{P},\pi_{e}}\left[Y_{t}-\sum_{a=1}^{K}1_{A_{t}=a}\theta^{\star}_{a}\right].$
Noting that $Y_{t}=Y_{t}(a)1_{A_{t}=a}$ and rearranging terms yields,
@@
To infer the average effect of treatment $j$ versus treatment $k$, we can then target $\eta^{T}\theta^{\star}$ for some contrast vector $\eta:=e_{j}-e_{k}$ (recall $e_{j}$ refers to the $j$-th standard basis vector in $\mathbb{R}^{K})$. For example $\eta=(-1,1)$ in the setting of $K=2$ would recover the average treatment effect $\mathbb{E}\left[Y(1)-Y(0)\right].$
@@
Similar to Zhang et al. (2021), one option is to consider estimators of the form
@@
When this is true conditionally, it will also be true marginally for any choice of policy. Therefore, under no model misspecification, we have at every time step $t$ that
@@
removing the need to consider the choice of policy when defining the target parameter. Assuming that the maximum occurs at a critical point of $m_{\theta}$, this implies that
@@
for all $X_{t},A_{t}$ and all $w_{t}\in\sigma(X_{t},A_{t},\mathcal{H}_{t-1})$. Thus, $w_{t}$ can be arbitrarily chosen based on $X_{t}$, $A_{t}$, and $\mathcal{H}_{t-1}$ without violating the first-order conditions. For this reason, previous work often uses $w_{t}$ as a free parameter to control the variance without having to worry that this type of re-weighting will change the target away from $\theta^{\star}$. As an example, Zhang et al. (2021) chooses $w_{t}=\sqrt{\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}(A_{t}|X_{t},\mathcal{H}_{t-1})}}$, which allows the variance of the score function to converge to a value independent of history. Other examples include online debiasing approaches (Deshpande et al., 2018; Khamaru et al., 2025) which use regularization to choose $w_{t}$ so that the variance of $\hat{\theta}_{T}$ is controlled.
@@
which is only true for the specific choice of $w_{t}=\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}(A_{t}|X_{t},\mathcal{H}_{t-1})}$. Thus, we are constrained to define an estimator such as
@@
Using inverse propensity weights ensures that $\hat{\theta}_{0}$ will be unbiased for $\theta^{\star}$, but we must now consider alternative methods to stabilize the variance of the estimator, which we will discuss in the next section.
@@
The estimator improves power when the model $f_{t}$ is a good estimator of $\mathbb{E}[Y_{t}|X_{t},A_{t}]$ by augmenting observations with predictions from a model. This approach is routinely used when estimating simple functionals like the average treatment effect in causal inference (Chernozhukov et al., 2018) and off-policy evaluation (Hadad et al., 2021) and is referred to as the augmented inverse probability weighted estimator in these settings. More recently, Zrnic and Candès (2024) introduces a similar modification of the score function of $M$-estimators with predictions derived from black box machine learning models, but limits their findings to the non-causal, non-adaptive observational setting.